Multiply: (1/7 – i)^2 I need help with this . Can someone please explain what to do for me using this example. I need to know the steps as well as how you got the answer. Thankyou for you time. Your are appreshated.
Answer by stanbon(75887)

(1/7 – i)^2
=(1/7 - i)(1/7 - i)
Use FOIL to get:
F:(1/7)^2 = 1/49
O:(1/7)(-)i= -1/7 i
I:(1/7)(-i) = -1/7 i
L:(-i)^2 = -1
-----------------------
Combine like terms to get:
(-48/49)-(2/7)i
========================
Cheers,
Stan H.